Listed below are a few events and transactions of Kodax Company.

Year 1

Jan. 2 Purchased 45,000 shares of Grecco Co. common stock for $451,000 cash. Grecco has 180,000 shares of common stock outstanding, and its activities will be significantly influenced by Kodax.
Sept. 1 Grecco declared and paid a cash dividend of $1.90 per share.
Dec. 31 Grecco announced that net income for the year is $492,900.


Year 2

June 1 Grecco declared and paid a cash dividend of $1.30 per share.
Dec. 31 Grecco announced that net income for the year is $710,400.
Dec. 31 Kodax sold 12,000 shares of Grecco for $171,500 cash.


Prepare journal entries to record the above transactions and events of Kodax Company. (Do not round intermediate calculations and
Answer #1
Debit Credit
Jan 2 Long-term investments—Grecco 451000
     Cash 451000
Sep 01 Cash 85500 =45000*1.90
     Long-term investments—Grecco 85500
Dec 31 Long-term investments—Grecco 123225 =492900*25%
        Earnings from long­-term investment 123225
Year 2
June 01 Cash 58500 =45000*1.30
     Long-term investments—Grecco 58500
Dec 31 Long-term investments—Grecco 177600 =710400*25%
        Earnings from long­-term investment 177600
Dec 31 Cash 171500
     Long-term investments—Grecco 162087 =607825/45000*12000
      Gain on sale of investment 9413
Workings:
Percentage of ownership 25% =45000/180000
Cost of investment 451000
Add: Share of net income 300825 =123225+177600
Less: Dividends received -144000 =85500+58500
Book value of Investment 607825
Answer #2

whta is the term for long term investment

answered by: Austin Rapp